### v1

@@ 3 Theoretical guarantees
Next, we want to show that $\boldsymbol{h}$ is continuous in $[0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$, and holomorphic in $\mathring{\mathcal{K}}$. First recall that affine maps are holomorphic and that composition of holomorphic maps is holomorphic [22]. The map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})$ is holomorphic by Assumption 2. For two generic matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, the map $\boldsymbol{B}\mapsto\boldsymbol{B}\odot\boldsymbol{A}$ is linear in $\boldsymbol{B}$, therefore it is holomorphic. Furthermore, the composition of two holomorphic functions is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)$ obtained from the composition of holomorphic maps is also holomorphic. Moreover, the map $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$ is linear in $\tilde{\boldsymbol{C}}$, so it is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$, which is the composition of $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(\boldsymbol{z})$ and $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$, is also holomorphic. Finally, the difference of two holomorphic functions is holomorphic. It then follows that $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$ is holomorphic. This shows that $h$ is holomorphic.
All that is left to verify is that the Lipschitz condition (18) holds for $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$. We start by observing that
$\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert$ $=\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})(\boldsymbol{u}-\boldsymbol{v})\rVert\leq\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\rVert\cdot\lVert\boldsymbol{u}-\boldsymbol{v}\rVert,$
@@
$\lVert\mathrm{row}_{i}(\tilde{\boldsymbol{M}}(t,\boldsymbol{z}))\rVert_{1}$ $=\sum_{j\in V}\lvert\tilde{M}_{ij}(t,\boldsymbol{z})\rvert$
$=\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})W(t)_{ij}-\delta_{ij}\sum_{k\in V}\tilde{C}_{ik}(t,\boldsymbol{z})W(t)_{ik}\rvert$
$=2\cdot\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert W(t)_{ij}$
$\leq 2\cdot\max_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert\cdot\sum_{j\in V}W(t)_{ij}<\infty.$
Since $\tilde{\boldsymbol{C}}$ is continuous in $(t,\boldsymbol{z})$ and $[0,T]\times\mathcal{K}$ is compact, the image $\tilde{\boldsymbol{C}}([0,T]\times\mathcal{K})$ is also compact. Hence, the maximum in the last line is finite. The sum in the last line is also finite, because $|V|$ is finite and $\boldsymbol{W}:[0,T]\mapsto\mathbb{R}^{|V|\times|V|}$ is continuous on a compact set. Therefore, $\lVert\boldsymbol{M}\rVert_{\infty}$ is bounded, and the Lipschitz condition (18) holds with $L=\lVert\boldsymbol{M}\rVert_{\infty\rightarrow\infty}<\infty$ for $h$.
Step 2: validity of the uniform bound (19)
We want to bound $\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}$. First, note that
(21) $\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{f}(\boldsymbol{z})\|_{2}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{u}}(T,\boldsymbol{z})\|_{2}.$
Now, let $\tilde{\boldsymbol{v}}(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\boldsymbol{v}(t,\boldsymbol{z})$, where $\boldsymbol{v}(t,\boldsymbol{z}):=\|\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\|_{2}^{2}=\langle\tilde{\boldsymbol{u}}(t,\boldsymbol{z}),\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\rangle$ and $\langle\boldsymbol{v},\boldsymbol{w}\rangle=\boldsymbol{v}^{*}\boldsymbol{w}$ for any $\boldsymbol{v},\boldsymbol{w}\in\mathbb{C}^{|V|}$. To make the notation lighter, we define $\dot{\tilde{\boldsymbol{u}}}(t,\boldsymbol{z})=\frac{\partial\tilde{\boldsymbol{u}}(t,\boldsymbol{z})}{\partial t}$. We have that:
$=\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle+\langle\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle$
$=\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle+\overline{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle}$
$=2\cdot\text{Re}\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle$
where $r(\tilde{\boldsymbol{M}})=\sup_{\boldsymbol{w}}\frac{|\langle\boldsymbol{w},\tilde{\boldsymbol{M}}\boldsymbol{w}\rangle|}{\langle\boldsymbol{w},\boldsymbol{w}\rangle}$ is the numerical radius of $\tilde{\boldsymbol{M}}$ (see, e.g., [24]). The last inequality follows from the relation
$\frac{\text{Re}\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle}{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle}\leq\frac{|\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle|}{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle}\leq r(\tilde{\boldsymbol{M}}).$
Moreover, by the Cauchy-Schwarz inequality,
(22) $r(\tilde{\boldsymbol{M}})=\sup_{\boldsymbol{w}}\frac{|\langle\boldsymbol{w},\tilde{\boldsymbol{M}}\boldsymbol{w}\rangle|}{\langle\boldsymbol{w},\boldsymbol{w}\rangle}\leq\sup_{\boldsymbol{w}}\frac{\|\tilde{\boldsymbol{M}}\boldsymbol{w}\|_{2}\|\boldsymbol{w}\|_{2}}{\|\boldsymbol{w}\|_{2}^{2}}=\|\boldsymbol{M}\|_{2\rightarrow 2},$
@@
Defining $\beta(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}$, we see that
$\frac{\partial\boldsymbol{v}(t,\boldsymbol{z})}{\partial t}\leq\beta(t)\boldsymbol{v}(t,\boldsymbol{z}),$
which, in turn, leads to
$\tilde{\boldsymbol{v}}^{\prime}(t)\leq\beta(t)\tilde{\boldsymbol{v}}(t).$
We can now apply Gronwall’s lemma [20], obtaining
@@
From now on, we will turn our attention back to scalar-valued functions $f_{v}:\mathbb{C}^{d}\mapsto\mathbb{C}$, as we wants to find a suitable sparse polynomial approximation to scalar-valued parametric map defined in (17) using the methods reviewed in Section 2.3. Note that analogous results could be derived in the vector-valued fields (and therefore, for the whole parametric solution of the diffusion map $\boldsymbol{y}\mapsto\boldsymbol{u}(\cdot,\boldsymbol{y})$) (see Remark 3.6).
3.1 Best $s$-term approximation rates
Based on Theorem 3.1, we can infer results about the convergence rate of the best $s$-term approximation error (recall Definition 1) for the solution map $f_{v}$ defined in equation (17). By combining [5, Theorem 3.6] and Theorem 3.1, we obtain the following corollary.
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$, be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $v\in V$ and $\boldsymbol{u}$ is the solution of problem (4). Moreover, assume that the compact set $\mathcal{K}\subseteq\mathbb{C}^{d}$ of assumption 2 is such that $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $B(u_{0},T,\boldsymbol{M)}$ be the bound in equation (20) and let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ be the sequence of coefficient of $f_{v}$ with respect to either the Chebyshev or Legendre basis, i.e., $f_{v}=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$. Then, $\forall s\in\mathbb{N}_{0}$, and $\forall p\in(0,q]$ the bound
$\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{B(u_{0},T,\boldsymbol{M})\cdot C(p,\boldsymbol{\rho},d)}{(s+1)^{\frac{1}{p}-\frac{1}{q}}}$
holds.
This result implies that the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ with respect to Legendre or Chebyshev polynomials decays algebraically fast in $s$. In particular, $f_{v}$ defined in equation (17) admits a holomorphic extension to some compact $\mathcal{K}$, therefore the best $s$-term approximation error is bounded from above as in equation (15). The term $\lVert f_{v}\rVert_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}$ in equation (15) is bounded as in Theorem 3.1, with $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$.
3.2 Convergence guarantees for least squares and compressed sensing
Theorem 3.1 will be now leveraged to show algebraic convergence rates for both least squares and compressed sensing (in particular, for weighted SR-LASSO decoders; recall (12)).
Corollary 2 (Convergence of least squares-based surrogates).
Let $10<\epsilon<$ 1, $\varrho$ be either the uniform or Chebyschev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independent and identically distributed drawn from the probability distribution $\varrho$. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution of equation (4) with the diffusivity coefficient $\boldsymbol{C}(t,\boldsymbol{y})$ satisfying Assumption 2 where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\{\psi_{\boldsymbol{\nu}}\}$ be either the Chebyshev or Legendre basis. Let $\mathcal{P}_{S}=\text{Span}(\{\psi_{\boldsymbol{\nu}}\})$. Then, there exists a set $S\subseteq\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f_{v}}=\sum_{\boldsymbol{\nu}\in S}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$ that solves the least squares problem (9) is unique and satisfies
The proof of Corollary 2 is based on [2, Theorem 6.1] and can be found in Appendix B.
Theorem 3.1 also implies uniform and nonuniform guarantees of algebraic convergence (using Theorems 7.12 and 7.13 in [5]) for compressed sensing. In particular, we aim at finding the minimizer of the weighted SR-LASSO decoder defined in (12).
Corollary 3 (Convergence of compressed sensing-based surrogates).
Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution to problem (4) with the diffusivity coefficient $C(t,\boldsymbol{y})$ satisfying Assumption 2, where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})$ be the sequence of coefficients of $f_{v}$ with respect to either the Chebyshev or Legendre basis, Let $m\in\mathbb{N}$ be the number of samples, drawing $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independently from the measure $\varrho$, and $10<\epsilon<$ 1.
Let
$\tilde{L}=\tilde{L}(m,d,\epsilon)=\kappa\cdot\log(2m)\cdot\big{[}\log(2m)\cdot\min\{\log(m)+d,\log(2d)\cdot\log(2m)\}+log(\epsilon^{-1})\big{]},$
with $\kappa>0$ a universal constant, $\tilde{m}=\tilde{m}(m,d,\epsilon)=m/\tilde{L}$ and $\Lambda=\Lambda_{n-1}^{HC}$ be the hyperbolic cross index set defined in (8) with $n=\lceil\tilde{m}\rceil$. Then the following holds with probability at least $1-\epsilon$. Any minimizer $\hat{\boldsymbol{c}}$ of (12) with the intrinsic weights $\boldsymbol{w}=(w_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\Lambda}$ where $w_{\boldsymbol{\nu}}=\|\psi_{\boldsymbol{\nu}}\|_{L^{\infty}(\mathcal{U})},\;\forall\boldsymbol{\nu}\in\Lambda$, and parameter $\lambda=1/(8\sqrt{\tilde{m}})$ satisfies, for every $0<p\leq 1$,
@@ 3.2 Convergence guarantees for least squares and compressed sensing
$\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq$ $\;\kappa_{1}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big{(}C_{1}+C_{2}\big{)}\cdot\tilde{m}^{1/2-1/p}+\kappa_{2}\cdot\|\boldsymbol{n}\|_{2}/\sqrt{m},$
(23) $\|f_{v}-\hat{f_{v}}\|_{L^{\infty}(\mathcal{U})}\leq$ $\;\kappa_{3}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big{(}C_{1}+C_{2}\big{)}\cdot\tilde{m}^{1-1/p}+\kappa_{4}\cdot\sqrt{\tilde{m}/m}\cdot\|\boldsymbol{n}\|_{2},$
[…]
Proof 3.3.
@@
The proof follow directly from [5, Theorem 7.12] and Theorem 3.1. Indeed, Theorem [5, Theorem 7.12] holds when [5, Assumption 2.3] holds. Now, we consider Theorem 3.1 with the following setting:
  • $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$,
  • $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component)
  • consequently, $\mathcal{O}=\mathring{\mathcal{K}}$.
Therefore, [5, Assumption 2.3] holds and so does [5, Theorem 7.12].

### v2

@@ 3 Theoretical guarantees
Next, we want to show that $\boldsymbol{h}$ is continuous in $0,T]^{2}\times\mathbb{C}^{|V|}\times\mathcal{K}$, and holomorphic in $\mathring{\mathcal{K}}$. First recall that affine maps are holomorphic and that composition of holomorphic maps is holomorphic [27]. The map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})$ is holomorphic by Assumption 2. For two generic matrices $\boldsymbol{A}$ and $\boldsymbol{B}$, the map $\boldsymbol{B}\mapsto\boldsymbol{B}\odot\boldsymbol{A}$ is linear in $\boldsymbol{B}$, therefore it is holomorphic. Furthermore, the composition of two holomorphic functions is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)$ obtained from the composition of holomorphic maps is also holomorphic. Moreover, the map $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$ is linear in $\tilde{\boldsymbol{C}}$, so it is holomorphic. Thus, the map $\boldsymbol{z}\mapsto\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$, which is the composition of $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(\boldsymbol{z})$ and $\tilde{\boldsymbol{C}}\mapsto\delta_{ij}\sum_{k}\tilde{C}_{ik}W_{ik}$, is also holomorphic. Finally, the difference of two holomorphic functions is holomorphic. It then follows that $\boldsymbol{z}\mapsto\tilde{\boldsymbol{C}}(t,\boldsymbol{z})\odot\boldsymbol{W}(t)-\boldsymbol{D}(\tilde{\boldsymbol{C}}(\boldsymbol{z}))$ is holomorphic. This shows that $h$ is holomorphic.
All that is left to verify is that the Lipschitz condition (20) holds for $\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})=\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\boldsymbol{u}$. We start by observing that $\lVert\boldsymbol{h}(t,s,\boldsymbol{u},\boldsymbol{z})-\boldsymbol{h}(t,s,\boldsymbol{v},\boldsymbol{z})\rVert=\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})(\boldsymbol{u}-\boldsymbol{v})\rVert\leq\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\rVert\cdot\lVert\boldsymbol{u}-\boldsymbol{v}\rVert$, where the norm in $\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\rVert$ is the matrix norm induced by $\|\cdot\|$. Here, we choose $\lVert\cdot\rVert=\lVert\cdot\rVert_{\infty}$. We now need to find an upper bound to $\lVert\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\rVert_{\infty\rightarrow\infty}=\max_{i\in V}\lVert\mathrm{row}_{i}(\tilde{\boldsymbol{M}}(t,\boldsymbol{z}))\rVert_{1}$. We know that
$\lVert\mathrm{row}_{i}(\tilde{\boldsymbol{M}}(t,\boldsymbol{z}))\rVert_{1}$ $=\sum_{j\in V}\lvert\tilde{M}_{ij}(t,\boldsymbol{z})\rvert=\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})W(t)_{ij}-\delta_{ij}\sum_{k\in V}\tilde{C}_{ik}(t,\boldsymbol{z})W(t)_{ik}\rvert$
$=2\cdot\sum_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert W(t)_{ij}\leq 2\cdot\max_{j\in V}\lvert\tilde{C}_{ij}(t,\boldsymbol{z})\rvert\cdot\sum_{j\in V}W(t)_{ij}<\infty.$
Since $\tilde{\boldsymbol{C}}$ is continuous in $(t,\boldsymbol{z})$ and $0,T]\times\mathcal{K}$ is compact, the image $\tilde{\boldsymbol{C}}([0,T]\times\mathcal{K})$ is also compact. Hence, the maximum in the last line is finite. The sum in the last line is also finite, because $|V|$ is finite and $\boldsymbol{W}:[0,T]\mapsto\mathbb{R}^{|V|\times|V|}$ is continuous on a compact set. Therefore, $\lVert\boldsymbol{M}\rVert_{\infty}$ is bounded, and the Lipschitz condition (20) holds with $L=\lVert\boldsymbol{M}\rVert_{\infty\rightarrow\infty}<\infty$.
Step 2: validity of the uniform bound (21)
We want to bound $\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}$. Note that
$\|\tilde{\boldsymbol{f}}\|_{L^{\infty}(\mathcal{H})}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{f}(\boldsymbol{z})\|_{2}=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{u}}(T,\boldsymbol{z})\|_{2}.$ (23)
Now, let $\tilde{\boldsymbol{v}}(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\boldsymbol{v}(t,\boldsymbol{z})$, where $\boldsymbol{v}(t,\boldsymbol{z}):=\|\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\|_{2}^{2}=\langle\tilde{\boldsymbol{u}}(t,\boldsymbol{z}),\tilde{\boldsymbol{u}}(t,\boldsymbol{z})\rangle$ and $\langle\boldsymbol{v},\boldsymbol{w}\rangle=\boldsymbol{v}^{*}\boldsymbol{w}$ for any $\boldsymbol{v},\boldsymbol{w}\in\mathbb{C}^{|V|}$. To make the notation lighter, we define $\dot{\tilde{\boldsymbol{u}}}(t,\boldsymbol{z})=\frac{\partial\tilde{\boldsymbol{u}}(t,\boldsymbol{z})}{\partial t}$. We have that
$=\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle+\langle\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}},\tilde{\boldsymbol{u}}\rangle=\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle+\overline{\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle}=2\cdot\text{Re}\langle\tilde{\boldsymbol{u}},\tilde{\boldsymbol{M}}\tilde{\boldsymbol{u}}\rangle$
@@
$r(\tilde{\boldsymbol{M}})=\sup_{\boldsymbol{w}}\frac{|\langle\boldsymbol{w},\tilde{\boldsymbol{M}}\boldsymbol{w}\rangle|}{\langle\boldsymbol{w},\boldsymbol{w}\rangle}\leq\sup_{\boldsymbol{w}}\frac{\|\tilde{\boldsymbol{M}}\boldsymbol{w}\|_{2}\|\boldsymbol{w}\|_{2}}{\|\boldsymbol{w}\|_{2}^{2}}=\|\tilde{\boldsymbol{M}}\|_{2\rightarrow 2},$ (24)
@@
Defining $\beta(t):=\sup_{\boldsymbol{z}\in\mathcal{H}}\|\tilde{\boldsymbol{M}}(t,\boldsymbol{z})\|_{2\to 2}$, we see that $\frac{\partial\boldsymbol{v}(t,\boldsymbol{z})}{\partial t}\leq\beta(t)\boldsymbol{v}(t,\boldsymbol{z})$, which, in turn, leads to $\tilde{\boldsymbol{v}}^{\prime}(t)\leq\beta(t)\tilde{\boldsymbol{v}}(t)$. We can now apply Gronwall’s lemma [29], obtaining
@@
which corresponds to (21). We conclude by showing that $\int_{0}^{T}\beta(t)dt<\infty$. Indeed, we have that $\beta$ is continuous in $t$, as $\|\cdot\|_{2\to 2}$ is continuous and $\tilde{\boldsymbol{M}}(t,\boldsymbol{z})$ is continuous. Therefore $\beta$ in bounded on $[0,T]$ and the integral is finite. This concludes the proof. ∎
From now on, we will turn our attention back to scalar-valued functions $f_{v}:\mathbb{C}^{d}\mapsto\mathbb{C}$, as we wants to find a suitable sparse polynomial approximation to scalar-valued parametric map defined in (19) using the methods reviewed in Section 2.3. Note that analogous results could be derived in the vector-valued fields (and therefore, for the whole parametric solution of the diffusion map $\boldsymbol{y}\mapsto\boldsymbol{u}(\cdot,\boldsymbol{y})$) (see Remark 3).
3.1 Best $s$-term approximation rates
Based on Theorem 1, we can infer results about the convergence rate of the best $s$-term approximation error (recall Definition 1) for the solution map $f_{v}$ defined in equation (19). By combining [1, Theorem 3.6] and Theorem 1, we obtain the following corollary.
Let $T>0$ and let the parametric diffusivity matrix $\boldsymbol{C}:[0,T]\times\mathbb{R}^{|V|\times|V|}$ satisfy Assumption 2. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$, be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $v\in V$ and $\boldsymbol{u}$ is the solution of problem (4). Moreover, assume that the compact set $\mathcal{K}\subseteq\mathbb{C}^{d}$ of assumption 2 is such that $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $B(u_{0},T,\boldsymbol{M)}$ be the bound in equation (22) and let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}$ be the sequence of coefficient of $f_{v}$ with respect to either the Chebyshev or Legendre basis, i.e., $f_{v}=\sum_{\boldsymbol{\nu}\in\mathbb{N}_{0}^{d}}c_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$. Then, $\forall s\in\mathbb{N}_{0}$, and $0<p\leq q<\infty$ it holds
$\sigma_{s}(\boldsymbol{c})_{q}\leq\frac{B(u_{0},T,\boldsymbol{M})\cdot C(p,\boldsymbol{\rho},d)}{(s+1)^{\frac{1}{p}-\frac{1}{q}}}.$
This result implies that the best $s$-term approximation error $\sigma_{s}(\boldsymbol{c})_{q}$ with respect to Legendre or Chebyshev polynomials decays algebraically fast in $s$. In particular, $f_{v}$ defined in equation (19) admits a holomorphic extension to some compact $\mathcal{K}$, therefore the best $s$-term approximation error is bounded from above as in equation (15). The term $\lVert f_{v}\rVert_{L^{\infty}(\mathcal{E}_{\boldsymbol{\rho}})}$ in equation (15) is bounded as in Theorem 1, with $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$.
3.2 Convergence guarantees for least squares and compressed sensing
Theorem 1 will be now leveraged to show algebraic convergence rates for both least squares and compressed sensing (in particular, for weighted SR-LASSO decoders; recall (12)).
Corollary 2 (Convergence of least squares-based surrogates).
Let $0<\epsilon<1$, $\varrho$ be either the uniform or Chebyshev measure on $\mathcal{U}=[-1,1]^{d}$, $m\geq 3$ and $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independent and identically distributed drawn from the probability distribution $\varrho$. Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution of equation (4) with the diffusivity coefficient $\boldsymbol{C}(t,\boldsymbol{y})$ satisfying Assumption 2 where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\{\psi_{\boldsymbol{\nu}}\}$ be either the Chebyshev or Legendre basis. Then, there exists a set $S\subset\mathbb{N}_{0}^{d}$ of cardinality $|S|\leq\lceil m/\log(m/\epsilon)\rceil$ such that the following holds with probability at least $1-\epsilon$. For any $\boldsymbol{n}\in\mathbb{C}^{m}$, the approximation $\hat{f_{v}}=\sum_{\boldsymbol{\nu}\in S}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$ that solves the least squares problem (9) is unique and satisfies, for every $0<p<1$,
The proof of Corollary 2, based on [26, Theorem 6.1], can be found in Appendix B.
Theorem 1 also implies uniform and nonuniform guarantees of algebraic convergence (using Theorems 7.12 and 7.13 in [1]) for compressed sensing. In particular, we aim at finding the minimizer of the weighted SR-LASSO decoder defined in (12).
Corollary 3 (Convergence of compressed sensing-based surrogates).
Let the map $f_{v}:\mathcal{U}=[-1,1]^{d}\to\mathbb{C}$ be defined by $f_{v}(\boldsymbol{y})=u_{v}(T,\boldsymbol{y})$, where $\boldsymbol{u}$ is the solution to problem (4) with the diffusivity coefficient $C(t,\boldsymbol{y})$ satisfying Assumption 2, where $\mathcal{E}_{\boldsymbol{\rho}}\subset\mathring{\mathcal{K}}$ for some $\boldsymbol{\rho}>\boldsymbol{1}$; let $\boldsymbol{c}=(c_{\boldsymbol{\nu}})$ be the sequence of coefficients of $f_{v}$ with respect to either the Chebyshev or Legendre basis, Let $m\in\mathbb{N}$ be the number of samples, drawing $\boldsymbol{y}_{1},\dots,\boldsymbol{y}_{m}$ independently from the measure $\varrho$, and $0<\epsilon<1$.
@@ 3.2 Convergence guarantees for least squares and compressed sensing
$\begin{split}\|f_{v}-\hat{f_{v}}\|_{L_{\varrho}^{2}(\mathcal{U})}\leq&\;\kappa_{1}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big(C_{1}+C_{2}\big)\cdot\tilde{m}^{1/2-1/p}+\kappa_{2}\cdot\|\boldsymbol{n}\|_{2}/\sqrt{m},\\ \|f_{v}-\hat{f_{v}}\|_{L^{\infty}(\mathcal{U})}\leq&\;\kappa_{3}\cdot B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})\cdot\big(C_{1}+C_{2}\big)\cdot\tilde{m}^{1-1/p}+\kappa_{4}\cdot\sqrt{\tilde{m}/m}\cdot\|\boldsymbol{n}\|_{2},\end{split}$
where $\kappa_{i}>0$, $i=1,\dots,4$ are universal constants, $\hat{f}=\sum_{\boldsymbol{\nu}\in\Lambda}\hat{c}_{\boldsymbol{\nu}}\psi_{\boldsymbol{\nu}}$, $B(u_{0},T,\boldsymbol{\rho},\boldsymbol{M})$ is the costant defined in (22), $C_{1}=C(\boldsymbol{\rho},p,d)$ is as in [1, Lemma 7.19] and $C_{2}=C(\boldsymbol{\rho},p,d)$ is as in [1, Lemma 3.13].
Proof.
@@
The proof follow directly from [1, Theorem 7.12] and Theorem 1. Indeed, Theorem [1, Theorem 7.12] holds when [1, Assumption 2.3] holds. Now, we consider Theorem 1 with the following setting: $\mathcal{H}=\mathcal{E}_{\boldsymbol{\rho}}$, $\mathcal{K}=\mathcal{E}_{\boldsymbol{\rho}^{\prime}}$, where $\boldsymbol{\rho}^{\prime}>\boldsymbol{\rho}$ (strict inequality applies to every component), and, consequently, $\mathcal{O}=\mathring{\mathcal{K}}$. Therefore, [1, Assumption 2.3] holds and so does [1, Theorem 7.12]. ∎